### v1

# Optimal navigation in two-dimensional regular and turbulent flows

Abstract
Zermelo’s navigation problem seeks the trajectory of minimal travel time between two points in a fluid flow. We address this problem for an agent – such as a micro-robot or active particle – that is advected by a two-dimensional flow, self-propels at a fixed speed smaller than or comparable to the characteristic flow velocity, and can steer its direction. The flows considered span increasing levels of complexity, from steady solid-body rotation to the Taylor-Green flow and fully developed turbulence in the inverse cascade regime. Although optimal control theory provides time-minimizing trajectories, these solutions become unstable in chaotic regimes realized for complex background flows. To design robust navigation strategies under such conditions, we apply reinforcement learning. Both action-value (Q-learning) and policy-gradient (one-step actor-critic) methods achieve successful navigation with comparable performance. Crucially, we show that agents trained on coarse-grained flows – retaining only large-scale features – generalize effectively to the full turbulent field. This robustness to incomplete flow information is essential for practical navigation in real-world oceanic and atmospheric environments.
I Introduction
The quest for the optimal path between two points, a cornerstone of navigation and control theory, finds profound expression in the classical Zermelo’s navigation problem. Originally formulated by Ernst Zermelo in his seminal work [zermelo1931navigationsproblem], the problem seeks the trajectory of minimal travel time for an agent moving with constant self-propulsion speed in a known fluid flow by adjusting the steering direction. Beyond its theoretical elegance, this framework holds critical importance for a variety of modern applications, from micro-robots navigating vascular networks for targeted drug delivery [nelson2010microrobots, park2017multifunctional, manzari2021targeted], to autonomous underwater and aerial vehicles operating in oceanic and atmospheric environments [trincavelli2008towards, lermusiaux2017future, bellemare2020autonomous, chai2020monitoring]. The core challenge lies in the interplay between the agent’s own propulsion and the advective power of the flow, which often dominates the dynamics, necessitating a control strategy that can exploit environmental forces rather than simply resist them.
The study of Zermelo’s problem has traditionally been approached through the lens of optimal control theory. Within this framework, one can derive a differential equation governing the evolution of the agent’s steering direction, which serves as the necessary condition for optimality (see, e.g., Ref. [bryson2018applied, Sec. 2.7]). For relatively simple flow fields, this equation admits analytical solutions [techy2011optimal], enabling the construction of optimal trajectories that can be used as benchmarks. In more complex environments, however, numerical techniques are required. The challenge is further complicated by the absence of a known initial condition for the steering equation, which in practice must be determined using shooting methods [kirk2004optimal].
This classical formulation faces two crucial limitations. First, calculating the optimal path requires perfect and global knowledge of the flow field. Second, the resulting optimal trajectories can be unstable. In systems exhibiting chaotic dynamics – characterized by exponential separation of nearby trajectories – infinitesimal uncertainties in initial conditions or control inputs are amplified, leading to drastic deviations from the planned route [angelo2009chaos, biferale2019zermelo]. In these cases, pre-computed optimal control strategies become ineffective, as the agent must perform additional actions to compensate deviations from the found path. Both limitations are violated in turbulent flows, which dominate oceanic and atmospheric environments [vallis2017atmospheric, boffetta2012two]. The requirement for perfect flow knowledge is unrealistic, as operational forecasts are inherently coarse-grained, forcing an agent to operate with incomplete information. Furthermore, turbulence is a canonical example of a chaotic system, meaning the stability condition is violated.
In response to these challenges, the field has explored alternative paradigms, including heuristic [piro2022optimal, monthiller2022surfing, calascibetta2023optimal] and data-driven methods [yoo2016path, colabrese2017flow, gustavsson2017finding, schneider2019optimal, biferale2019zermelo, buzzicotti2020optimal, alageshan2020machine, gunnarson2021learning, nasiri2023optimal]. The rise of machine learning has opened new avenues, with reinforcement learning (RL) emerging as a particularly powerful model-free framework for complex control [sutton1998reinforcement]. The RL formulation treats navigation as a sequential decision-making process in which an agent learns a policy – a mapping from states (e.g., position) to actions (e.g., heading adjustments) – by maximizing a cumulative reward, typically defined in terms of minimizing travel time to a target. Early studies have highlighted the promise of this approach. For instance, the work of Colabrese et al. [colabrese2017flow] demonstrated that smart active particles could use Q-learning to navigate vortex flows, effectively discovering physical strategies like surfing on specific flow structures. Similarly, Biferale et al. [biferale2019zermelo] explored the use of RL to solve Zermelo’s problem in synthetic two-dimensional turbulent flows, showing that agents can learn to leverage flow structures to minimize travel time. However, a systematic investigation across a hierarchy of flows from regular to fully turbulent, and crucially, an assessment of the robustness and generalization capabilities of learned policies under realistic information constraints, remains an open area of research.
This work directly addresses this gap by presenting a comprehensive study of optimal navigation across a spectrum of two-dimensional flows. We examine environments of increasing complexity, from steady solid-body rotation to the advection by Taylor-Green vortices and, finally, to turbulent flows in the inverse cascade regime. The focus on two-dimensional flows is motivated by both computational tractability and their relevance to large-scale atmospheric and oceanic currents [vallis2017atmospheric, boffetta2012two]. Our analysis first confirms that while optimal control theory yields nominal time-minimizing trajectories, its practical utility is significantly reduced in chaotic regimes due to the instability discussed above. To develop more robust strategies, we then turn to RL and examine two algorithmic classes: an action-value approach based on Q-learning and a policy-gradient approach, implemented as a one-step actor-critic algorithm [sutton1998reinforcement]. We compare the resulting RL-derived trajectories with those obtained through classical optimal control to assess the capabilities of RL algorithms. The central and most significant finding of our study is that RL agents trained not on a full turbulent field, but on a coarse-grained version retaining only large-scale flow features, demonstrate remarkable generalization by successfully navigating within the fully resolved turbulent environment. This ability to operate effectively with incomplete, low-resolution information is not only of theoretical interest but also a critical step toward the practical deployment of autonomous systems in real-world oceanic and atmospheric settings.
By maintaining a detailed presentation, our work serves a dual purpose: advancing the research frontier and providing a valuable reference for educational purposes. The code that supports the findings of this study are openly available at https://github.com/parfenyev/opt-navig. The remainder of this paper is organized as follows. In Section LABEL:sec:2, we formulate the Zermelo’s problem and introduce the equations of motion for the agent. Section LABEL:sec:3 describe the hierarchy of flows we study. Section LABEL:sec:4 details the application of optimal control theory and analyzes the stability of the resulting trajectories. Section LABEL:sec:5 explains our reinforcement learning framework, including the state and action spaces, reward structure, and the implementation details of the Q-learning and one-step actor-critic algorithms. Section LABEL:sec:6 is devoted to navigation in turbulent environment with incomplete flow information. Finally, Section LABEL:sec:7 summarizes our findings.
II Problem Setup
[…]
sec:2
We study the motion of an agent navigating a two-dimensional flow field $\bm{v}(\bm{r},t)$, which is governed by the kinematic equation
[…]
where $\bm{R}_{t}=(X_{t},Y_{t})$ is the position of the agent at time $t$ in Cartesian coordinates. The dynamics combines passive advection by the background flow $\bm{v}(\bm{R}_{t},t)$, which represents the external environmental velocity such as oceanic or atmospheric currents, and active self-propulsion at constant speed $V_{0}$. The steering direction $\theta_{t}$ serves as the control variable that the agent can adjust in real-time. The objective of Zermelo’s problem [zermelo1931navigationsproblem] is to find a control policy $\theta_{t}=\theta(\bm{R}_{t},t)$ that minimizes the total travel time $T$ for a journey between two predefined points $\bm{r}_{A}$ and $\bm{r}_{B}$. This challenge originated in the context of airship flight [ebbinghaus2007ernst] and remains a canonical problem in optimal control theory.
The nature of the solution is dependent on the relative strength of the flow. In the regime, where the agent’s speed $V_{0}$ significantly exceeds the characteristic flow velocity, one may expect that a trivial strategy of continuously orienting the propulsion vector directly toward the target $\bm{r}_{B}$ is close to optimal. The problem becomes more challenging in the opposing regime, where the flow exerts a strong influence. Under these conditions, the agent must strategically ride the current, often making counter-intuitive moves away from the target to exploit favorable flow structures for an overall reduction in travel time. This latter complex regime is the primary focus of our work.
III Background Flows
[…]
sec:3
This section introduces the hierarchy of background flows considered in our study. The first two examples are adopted from Ref. [techy2011optimal], and for them the Zermelo’s problem admits a semi-analytical solution. We then move on to more complex flows designed to test the performance of our methods in increasingly challenging environments.
III.1 Steady vortex flow
[…]
sec:Ex_A
In the first example, the background velocity field corresponds to steady solid-body rotation and it is given by
$\bm{v}(x,y)=\left(-\omega y,\;\omega x\right)^{T}.$ (2)
The starting point is $(x_{A},y_{A})=\left(1/2,\;\sqrt{3}/2\right)$ and our aim is to reach the end point $(x_{B},y_{B})=(1,\;0)$ in the shortest time. The agent’s speed $V_{0}=1$, and we set $\omega=0.9$. Note that traveling upstream along a streamline is not an optimal solution.
III.2 Steady sink and time-varying vortex flow
[…]
sec:Ex_B
In the second example, we consider a divergent and time-varying flow. The background velocity is given by
[…]
The flow model represents a constant sink plus a vortex that linearly changes vorticity with time. The goal is to start from the point $(x_{A},y_{A})=\left(1/2,\;\sqrt{3}/2\right)$ and reach the point $(x_{B},y_{B})=(1,\;0)$ in the shortest time. The agent’s speed is $V_{0}=1$.
III.3 Taylor-Green vortices
[…]
sec:Ex_C
In the third example, the background velocity is incompressible, $\mathrm{div}\;\bm{v}=0$, and it is given by
$label{eq:TG-flow}\bm{v}(x,y)=\begin{pmatrix}u_{0}\sin(kx)\cos(ky)\\ -u_{0}\cos(kx)\sin(ky)\end{pmatrix}$. (4)
The corresponding vorticity is $\varpi=\partial_{x}v_{y}-\partial_{y}v_{x}=2u_{0}k\sin(kx)\sin(ky)$. Thus, the velocity field looks like a checkerboard of counter-rotating vortices. A similar velocity field can be generated on the fluid surface due to the nonlinear interaction of two orthogonal standing waves [filatov2016nonlinear, filatov2016generation, francois2017wave, xia2019generation, abella2020measurement].
We assume that the motion is confined in space within $0,L]^{2}$ domain, where $L=2\pi$. We set $u_{0}=1$, $k=3$, and the agent’s speed is $V_{0}=0.1$. The starting point is $(x_{A},y_{A})=\left(2\pi/3,\;\pi/3\right)$ and the end point is $(x_{B},y_{B})=(3\pi/2,\;3\pi/2)$. The main challenge in this example is to escape vortex traps along the way.
III.4 2D Turbulence
[…]
sec:Ex_D
Finally, we consider a two-dimensional turbulent flow, which is generated by numerically solving the incompressible forced 2D Navier-Stokes equations for a fluid of unit density
$label{eq:NS-equation}\partial_{t}\bm{v}+(\bm{v}\cdot\nabla)\bm{v}=-\nabla p+\nu\nabla^{2}\bm{v}-\alpha\bm{v}+\bm{f},\quad\nabla\cdot\bm{v}=0,$ (5)
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\nu=5\cdot 10^{-3}$ is the kinematic viscosity, $\alpha=0.2$ is the linear bottom drag, and $\bm{f}$ is an external random forcing. The equations are solved in a doubly periodic square domain of size $L=2\pi$. The forcing $\bm{f}$ is isotropic in space and shortly correlated in time, with a Gaussian spatial spectrum centered at wavenumber $k_{f}=5.5$ and having a standard deviation $\delta_{f}=1.0$. The energy injection rate $\epsilon=\langle\bm{v}\cdot\bm{f}\rangle$ is an externally controlled parameter set to $\epsilon=10$.
DNS results are obtained by integrating (LABEL:eq:NS-equation) using the GeophysicalFlows.jl pseudospectral code [GeophysicalFlowsJOSS], fully dealiazed by the two-thirds rule. We use a $512^{2}$ grid with a fixed integration time step $dt=2\cdot 10^{-4}$, resulting in a Courant number <0.3<0.3. The forcing correlation time is also $dt$. Starting from rest, the simulation is run until it reaches a non-equilibrium stationary state, observed by the saturation of the total kinetic energy. Snapshots of the velocity field are then output at an interval of $\tau=5\cdot 10^{-3}$ over a total duration of $T_{tot}=6.15$, yielding $1230$ snapshots for analysis.
The resulting flow is characterized by dimensionless numbers $Re=UL/\nu\approx 8\cdot 10^{3}$ and $Rh=U/(\alpha L)\approx 5$, where $U\approx 6.3$ is the root-mean-square velocity. The mean energy spectrum, shown in Fig. LABEL:fig:1, confirms the established phenomenology of forced 2D turbulence. In the inverse energy cascade range $k<k_{f}$, the spectrum follows the theoretical $E(k)\propto k^{-5/3}$ scaling [boffetta2012two], while in the direct enstrophy cascade range $k>k_{f}$, the bottom drag leads to a steeper slope than the classical $E(k)\propto k^{-3}$ prediction [valadao2025spectrum].
To evaluate our navigation algorithms, we employ both instantaneous snapshot and the full time-dependent flow data. The agent’s self-propulsion speed is set to $V_{0}\approx 4.5$, which corresponds to $30\%$ of the maximum flow velocity. Experimentally, flows with similar statistical properties can be excited by Faraday waves on the fluid surface [von2011double, francois2013inverse, francois2014three, colombi2021three, colombi2022coexistence] or by the Lorentz force in a thin conducting fluid layer over a magnet array [sommeria1986experimental, paret1998intermittency, xia2009spectrally, bardoczi2014experimental, orlov2018large, zhu2024flow].
[…]
IV Optimal Control Theory
[…]
sec:4
In this section, we describe how to apply control theory to determine the time-minimizing trajectories for an agent moving from point $\bm{r}_{A}$ to point $\bm{r}_{B}$. First, we derive a necessary condition that the steering direction $\theta_{t}$ must satisfy along an optimal trajectory. Our analysis assumes only a basic familiarity with the calculus of variations and draws heavily on Ref. [bryson2018applied, Sec. 2.7]. For readers acquainted with Pontryagin’s minimum principle, a similar result is derived in Ref. [techy2011optimal]. Following the derivation of the Zermelo equation, we apply it to the specific examples outlined in the previous section.
IV.1 Zermelo equation
The problem of finding the minimum-time trajectory and the corresponding travel time $T$ can be addressed through variational calculus. We seek to minimize the functional $\int_{0}^{T}dt\;L(t,\bm{R}_{t},\dot{\bm{R}}_{t})$, where $\bm{R}_{t}$ obeys the equation of motion (LABEL:eq:motion) and boundary conditions $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. This constitutes a constrained optimization problem that can be solved using the method of Lagrange multipliers. For the minimum-time problem, the cost function is simply $L(t,\bm{R}_{t},\dot{\bm{R}}_{t})\equiv 1$, yielding the extended functional
$\begin{split}J=\int\limits_{0}^{T}dt\Big(1-\lambda_{X}(t)\left[V_{0}\cos\theta_{t}+v_{x}(t,X_{t},Y_{t})-\dot{X}_{t}\right]\\ -\lambda_{Y}(t)\left[V_{0}\sin\theta_{t}+v_{y}(t,X_{t},Y_{t})-\dot{Y}_{t}\right]\Big).\end{split}$ (6)
The Lagrange multipliers $\bm{\lambda}_{t}=(\lambda_{X}(t),\lambda_{Y}(t))$ play a role analogous to canonical momentum in classical mechanics. Following this analogy, we introduce the Hamiltonian $H=-1+\lambda_{X}(t)\left[V_{0}\cos\theta_{t}+v_{x}(t,X_{t},Y_{t})\right]+\lambda_{Y}(t)\left[V_{0}\sin\theta_{t}+v_{y}(t,X_{t},Y_{t})\right]$, and then the functional $J$ can be compactly written as
$J=\int_{0}^{T}dt\left(\bm{\lambda}_{t}\cdot\dot{\bm{R}}_{t}-H\right).$ (7)
Next, the necessary conditions for an extremum are derived by setting the first variation of the functional to zero, $\delta J=0$. However, in contrast to classical mechanics, the terminal time $T$ is not specified, and we should take into account its differential changes:
$\begin{split}\delta J=\int_{0}^{T}dt\Big[-\dfrac{\partial H}{\partial\bm{R}_{t}}\delta\bm{R}_{t}-\dfrac{\partial H}{\partial\bm{\lambda}_{t}}\delta\bm{\lambda}_{t}-\dfrac{\partial H}{\partial\theta_{t}}\delta\theta_{t}\\ +\dot{\bm{R}}_{t}\cdot\delta\bm{\lambda}_{t}+\bm{\lambda}_{t}\cdot\delta\dot{\bm{R}}_{t}\Big]+\delta T(\bm{\lambda}_{T}\cdot\dot{\bm{R}}_{T}-H(T)).\end{split}$ (8)
The last term in the square brackets can be integrated by parts, which produces additional boundary term $\bm{\lambda}_{T}\cdot\delta\bm{R}_{T}$. As the position of the final point is fixed, we get $d\bm{R}_{T}=\delta\bm{R}_{T}+\dot{\bm{R}}_{T}\delta T=0$, and therefore
$\begin{split}\delta J=\int_{0}^{T}dt\Bigg[\left(\dot{\bm{R}}_{t}-\dfrac{\partial H}{\partial\bm{\lambda}_{t}}\right)\delta\bm{\lambda}_{t}-\dfrac{\partial H}{\partial\theta_{t}}\delta\theta_{t}\\ -\left(\dfrac{\partial H}{\partial\bm{R}_{t}}+\dot{\bm{\lambda}}_{t}\right)\delta\bm{R}_{t}\Bigg]-H(T)\delta T.\end{split}$ (9)
For $\delta J=0$ to hold for arbitrary variations $\delta\bm{R}_{t}$, $\delta\bm{\lambda}_{t}$, $\delta\theta_{t}$, and $\delta T$, their coefficients must vanish identically. This yields the following necessary conditions for the optimal path:
$\begin{split}\dot{\bm{R}}_{t}=\dfrac{\partial H}{\partial\bm{\lambda}_{t}},\;\dot{\bm{\lambda}}_{t}=-\dfrac{\partial H}{\partial\bm{R}_{t}},\;\dfrac{\partial H}{\partial\theta_{t}}=0,\;\bm{\lambda}_{T}\cdot\dot{\bm{R}}_{T}=1.\end{split}$ (10)
The obtained system is governed by four differential equations (for $\bm{R}_{t}$ and $\bm{\lambda}_{t}$) with four boundary conditions (the fixed start and end points, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$), one optimality condition for the control $\theta_{t}$, and one transversality condition $\bm{\lambda}_{T}\cdot\dot{\bm{R}}_{T}=1$ for the free terminal time $T$. The state equation $\dot{\bm{R}}_{t}=\partial H/\partial\bm{\lambda}_{t}$ simply restates the original equations of motion (LABEL:eq:motion). From $\dot{\bm{\lambda}}_{t}=-\partial H/\partial\bm{R}_{t}$ and $\partial H/\partial\theta_{t}=0$, we obtain
$\dot{\lambda}_{t}^{x}=-\lambda_{t}^{x}\partial_{x}v_{x}(t,\bm{R}_{t})-\lambda_{t}^{y}\partial_{x}v_{y}(t,\bm{R}_{t}),$ (11)
$\dot{\lambda}_{t}^{y}=-\lambda_{t}^{x}\partial_{y}v_{x}(t,\bm{R}_{t})-\lambda_{t}^{y}\partial_{y}v_{y}(t,\bm{R}_{t}),$ (12)
$\tan\theta_{t}=\lambda_{t}^{y}/\lambda_{t}^{x}.$ (13)
The optimal control is determined by $\theta_{t}=\arctan(\lambda_{t}^{y}/\lambda_{t}^{x})$. A more useful form is the Zermelo equation, derived by differentiating this expression with respect to time
[…]
where the velocity gradients are evaluated along the agent trajectory $\bm{R}_{t}$. This equation describes how the vehicle must continuously adjust its heading in response to the spatial gradients of the flow field to maintain a minimum-time trajectory.
The complete Two-Point Boundary Value Problem (TPBVP) is thus defined by coupling the equations of motion (LABEL:eq:motion) with the heading evolution equation (LABEL:eq:angle). The boundary conditions are the fixed start and end positions, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. The initial heading $\theta_{0}$ is not specified, making this a free condition. The standard numerical approach to solve this TPBVP is the shooting method: an initial guess for $\theta_{0}$ is iteratively refined until the resulting trajectory, obtained by forward integration of the system, satisfies the terminal condition $\bm{R}_{T}=\bm{r}_{B}$ within a desired tolerance. More details on the numerical procedure will be provided later when discussing specific examples. Note that once the initial heading $\theta_{0}$ is known, the optimal path can be integrated autonomously, provided the agent can measure the local velocity field gradients, see Eq. (LABEL:eq:angle).
IV.2 Example 1: Steady vortex flow
[…]
sec:OPC_A
For the background velocity field $\bm{v}(x,y)=\left(-\omega y,\;\omega x\right)^{T}$, the Zermelo equation (LABEL:eq:angle) has a simple solution $\theta_{t}=\theta_{0}+\omega t$. Integrating the equations of motion (LABEL:eq:motion) from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$ yields the trajectory:
$X_{t}=$ $x_{A}\cos(\omega t)+V_{0}t\cos(\omega t)\cos\theta_{0}$
$-y_{A}\sin(\omega t)-V_{0}t\sin(\omega t)\sin\theta_{0},$
$Y_{t}=$ $y_{A}\cos(\omega t)+V_{0}t\cos(\omega t)\sin\theta_{0}$
$+x_{A}\sin(\omega t)+V_{0}t\sin(\omega t)\cos\theta_{0}.$
The initial heading $\theta_{0}$ and total travel time $T$ are determined by enforcing the terminal condition $(X_{T},Y_{T})=(x_{B},y_{B})$. For the numerical parameters specified in Section LABEL:sec:Ex_A, we obtain:
$\theta_{0}\approx-1.93592,\quad T\approx 1.97494.$ (17)
The found optimal trajectory is illustrated in Fig. LABEL:fig:2 by a white line. The figure also includes trajectories corresponding to small deviations in the initial heading $\theta_{0}$ and slight perturbations in the initial position $\bm{r}_{A}$. Overall, these trajectories stay close to the optimal path and successfully reach a narrow neighborhood of the target point. Based on expressions (IV.2) and (IV.2), it can be concluded that the initial perturbations exhibit linear growth with time rather than exponential growth, indicating the stability of the optimal trajectory.
[…]
IV.3 Example 2: Steady sink and time-varying vortex flow
[…]
sec:OPC_B
For the background velocity field (LABEL:eq:sink+rotation), the Zermelo equation (LABEL:eq:angle) has a solution $\theta_{t}=\theta_{0}+t/2-t^{2}/2$. Then, the equations of motion (LABEL:eq:motion) can be integrated from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$
$X_{t}=$ $\dfrac{e^{-0.3t}}{3}\Bigg[3x_{A}\cos\left(\dfrac{t(t-1)}{2}\right)+3y_{A}\sin\left(\dfrac{t(t-1)}{2}\right)$ (18)
$+10V_{0}(e^{0.3t}-1)\cos\left(\dfrac{t(t-1)}{2}-\theta_{0}\right)\Bigg],$
$Y_{t}=$ $\dfrac{e^{-0.3t}}{3}\Bigg[3y_{A}\cos\left(\dfrac{t(t-1)}{2}\right)-3x_{A}\sin\left(\dfrac{t(t-1)}{2}\right)$ (19)
$-10V_{0}(e^{0.3t}-1)\sin\left(\dfrac{t(t-1)}{2}-\theta_{0}\right)\Bigg].$
The found trajectory should also pass through the terminal point $(X_{T},Y_{T})=(x_{B},y_{B})$, which allows us to obtain the initial heading $\theta_{0}$ and travel time $T$. For the numerical parameters specified in Section LABEL:sec:Ex_B, we obtain:
$\theta_{0}\approx-0.77543,\quad T\approx 1.03008.$ (20)
Figure LABEL:fig:3 presents the optimal trajectory (white line) superimposed on the velocity field at the final time $T$. The figure also illustrate the system’s sensitivity by displaying trajectories with small errors in the initial heading $\theta_{0}$ and initial position $\bm{r}_{A}$. All trajectories remain near the optimal path and successfully approach the destination point. The complete time-dependent evolution is provided in the video, see Ref. [repository]. Based on expressions (18) and (19), it can be concluded that the initial small perturbations $\delta X_{0},\delta Y_{0},\delta\theta_{0}$ do not exhibit exponential growth with time, indicating the stability of the optimal trajectory.
[…]
IV.4 Example 3: Taylor-Green vortices
[…]
sec:OPC_C
Next, we consider the Taylor-Green flow, see Section LABEL:sec:Ex_C. Now, the optimal control equations must be solved numerically. To define the problem over a fixed integration interval, it is convenient to introduce a rescaled time variable $\tau=t/T$, such that $\tau\in[0,1]$. Using this transformation, the governing equations take the form
$\dot{X}_{\tau}=u_{0}T\sin(kX_{\tau})\cos(kY_{\tau})+V_{0}T\cos(\theta_{\tau}),$ (21)
$\dot{Y}_{\tau}=-u_{0}T\cos(kX_{\tau})\sin(kY_{\tau})+V_{0}T\sin(\theta_{\tau}),$ (22)
$\dot{\theta}_{\tau}=u_{0}kT\big[\cos(kX_{\tau})\cos(kY_{\tau})\sin(2\theta_{\tau})$ (23)
$+\sin(kX_{\tau})\sin(kY_{\tau})\big],$
with the boundary conditions $(X_{0},Y_{0})=(x_{A},y_{A})$ and $(X_{1},Y_{1})=(x_{B},y_{B})$.
The system of equations can be numerically integrated starting from the initial position $\bm{r}_{A}$ for a prescribed initial steering angle $\theta_{0}$ and total travel time $T$. However, these two parameters are unknown and must be adjusted so that the final position $(X_{1},Y_{1})$ matches the desired target point $\bm{r}_{B}$. To achieve this, we employ a shooting method followed by fine-tuning optimization. Initially, approximate values of $\theta_{0}$ and $T$ are chosen to bring the trajectory close to the target. These parameters are then refined to minimize the discrepancy
[…]
between the simulated and desired endpoints. For optimization, we use gradient- and Hessian-based methods. In particular, our implementation first applies the Adam optimizer [kingma2014adam], followed by the BFGS algorithm [nocedal2006numerical]. Efficient computation of the loss gradients $\partial{\mathcal{L}}/\partial\theta_{0}$ and $\partial{\mathcal{L}}/\partial\theta_{0}$ requires the forward integration scheme to support automatic differentiation [baydin2018automatic]. Moreover, the integration must be performed with high numerical accuracy. Convergence tests with progressively reduced tolerances are essential to ensure that the solution is independent of the integration precision.
[…]
For the numerical parameters specified in Section LABEL:sec:Ex_C, we obtain:
$label{eq:opt-control-taylor-green}\theta_{0}\approx 1.56079,\quad T\approx 12.2038.$ (25)
The corresponding optimal trajectory is shown in Fig. LABEL:fig:4 by a white line. The figure also illustrate the system’s sensitivity by depicting trajectories perturbed by small deviations in the initial heading $\theta_{0}\pm\pi/9000$ and by displacements $\Delta=10^{-3}$ in the initial $x_{A}$ and $y_{A}$ positions. Although these perturbations are considerably smaller than those examined in previous examples, the resulting trajectories fail to reach the target $\bm{r}_{B}$.
To make a more quantitative stability analysis of the optimal trajectory, we consider the evolution of infinitesimal perturbations,
$(\delta X_{t},\delta Y_{t},\delta\theta_{t})^{T}=\hat{W}(t)(\delta X_{0},\delta Y_{0},\delta\theta_{0})^{T},$ (26)
where $\hat{W}(t)$ is the linearized evolution matrix that can be obtained numerically. A polar decomposition $\hat{W}(t)=\hat{V}(t)\hat{R}(t)$ allows us to separate the effects of rotation, represented by the orthogonal matrix $\hat{R}(t)$, from those of stretching, characterized by the positive-definite matrix $\hat{V}(t)$. The eigenvalues $h_{i}(t)$ of $\hat{V}(t)$ (equal to singular values of $\hat{W}(t)$) quantify the magnitude of stretching along different directions in phase space and are used to compute the finite-time Lyapunov exponents (FTLEs), defined as
$\lambda_{i}(t)=\dfrac{1}{t}\ln h_{i}(t),\quad i=1,2,3.$ (27)
The corresponding results are presented in Fig. LABEL:fig:5. The presence of a positive maximal FTLE indicates that the optimal trajectory is unstable with respect to small perturbations on the time scales needed to reach the target. This instability limits the practical applicability of the pre-computed optimal control strategy, as minor deviations in initial conditions can lead to significant trajectory divergence. Moreover, the resulting sensitivity makes it difficult to find suitable initial guess for parameters $\theta_{0}$ and $T$, that lead to a successful solution of the optimization problem and finding optimal trajectories that reach the target. Despite these challenges, computing optimal trajectories can still be practically valuable if the agent additionally employs heuristic strategies to overcome noise and remain near the optimal path [piro2022optimal]. A detailed discussion of such strategies lies beyond the scope of this work.
[…]
IV.5 Example 4: Snapshot of 2D Turbulence
[…]
sec:OPC_D
Next, we apply the numerical scheme described above to solve the optimal control equations for the steady background flow corresponding to a snapshot of two-dimensional turbulence, see Section LABEL:sec:Ex_D. The start and end points are chosen as $\bm{r}_{A}=(1.2,5.5)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Since the optimal trajectory only satisfies the local extremum condition, the resulting equations may, in general, admit multiple solutions. To illustrate this, Fig. LABEL:fig:6 presents two distinct optimal trajectories (shown as white lines) corresponding to different initial heading angles $\theta_{0}$ and total travel times $T$:
$\theta_{0}\approx-0.883601,\quad T\approx 0.819998,$ (28)
$\theta_{0}\approx-0.863081,\quad T\approx 1.250000.$ (29)
Along the first (lower) trajectory, the agent exploits two vortices encountered on its path to increase speed and reach the target more quickly. In contrast, along the second (upper) trajectory, the agent first uses a nearby vortex to enter a region of weak external flow, where it can maneuver efficiently even with a modest self-propulsion speed $V_{0}$, and then leverages a subsequent current to approach the target. Both trajectories demonstrate effective use of the surrounding flow field, though the first achieves a faster overall transit.
[…]
Figure LABEL:fig:6 also illustrates the sensitivity of the faster trajectory to small perturbations in the initial control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minute variations in either of these quantities lead to trajectories that deviate significantly from the optimal path and ultimately fail to reach the target. This observation is further supported by the analysis of FTLEs presented in Fig. LABEL:fig:7, where the maximal FTLE attains positive values, confirming that the corresponding trajectory is unstable on the considered time scales. The observed sensitivity to small perturbations reflects an instability in the optimal control dynamics, associated with the agent’s self-propulsion activity [biferale2019zermelo].
To better understand this phenomenon, it is instructive to contrast it with the behavior of passive tracers in a steady two-dimensional incompressible flow. For such passive particles, motion occurs strictly along the streamlines, which constrains their trajectories to be regular and precludes the onset of chaotic dynamics. In the present case, however, the agent possesses self-propulsion and an additional degree of freedom associated with the steering direction $\theta_{t}$, which breaks the integrability of the purely passive system. As a result, the considered active system becomes far more susceptible to chaotic behavior, with small perturbations in initial conditions potentially leading to large deviations in the resulting trajectories.
[…]
IV.6 Example 5: 2D Turbulence
[…]
sec:OPC_E
In our final example, we consider a background velocity field corresponding to a fully time-dependent two-dimensional turbulent flow, see Section LABEL:sec:Ex_D. The starting and ending points are set to $\bm{r}_{A}=(1.4,3.0)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Figure LABEL:fig:8 presents the optimal trajectory (depicted as a white line), which corresponds to the initial heading angle $\theta_{0}$ and total travel time $T$ given by
$\theta_{0}\approx-0.064968,\quad T\approx 0.700001,$ (30)
superimposed on the velocity field at the final time $T$ (a video demonstration is available in Ref. [repository]).
[…]
This trajectory represents one of potentially many solutions to the optimal control equations, reflecting the existence of multiple local extrema in the optimization landscape. While we cannot prove that the presented solution represents the absolute minimum travel time (i.e., the global optimum), it has been selected for illustration since it yields the shortest travel time among all optimal trajectories identified in our study. As in the previous example, Fig. LABEL:fig:8 highlights the sensitivity of the optimal trajectory to small perturbations in the initial conditions – both in the control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minor deviations in these parameters lead to trajectories that fail to reach the target, emphasizing the instability and complex structure of the underlying dynamical system. The positive value of the maximal FTLE, shown in Fig. LABEL:fig:9, further confirms the trajectory’s instability over the characteristic travel-time scale.
IV.7 Summary
The optimal control framework provides a systematic method for determining time-minimizing trajectories between two points in a fluid flow for an active agent. In simple background flows, these trajectories are robust to small perturbations in initial position and steering direction, making them suitable for practical navigation. Once the initial steering angle $\theta_{0}$ is set, the agent can autonomously adjust its heading according to Eq. (LABEL:eq:angle), provided it can measure the local velocity field gradients. In more complex flows, finding optimal trajectories becomes considerably more difficult. The resulting paths often exhibit sensitivity to initial conditions and become unstable, limiting the direct application of a naive optimal control strategy in realistic environments. Nevertheless, such trajectories remain valuable. First, the optimal paths reveal how an agent can exploit the structure of the ambient flow to minimize travel time, thereby providing fundamental insight into efficient locomotion strategies. When supplemented with heuristic mechanisms designed to mitigate the influence of environmental noise and modeling uncertainties, these theoretically optimal solutions can still guide the construction of robust navigation algorithms [piro2022optimal]. Second, the optimal trajectories serve as valuable benchmarks: they establish theoretical lower bounds on travel time and offer reference solutions for evaluating the performance of alternative, stability-oriented navigation methods.
[…]
V Reinforcement Learning
[…]
sec:5
As discussed above, the optimal control approach can be highly sensitive to small perturbations, which limits its robustness in complex flow environments. To overcome this limitation, in this section we employ reinforcement learning (RL) to develop stable navigation strategies. Specifically, we consider two classes of RL algorithms: an action-value method based on Q-learning and a policy-gradient method implemented through a one-step actor-critic framework. We briefly outline the principles of these methods and then apply them to various background flow scenarios, comparing their performance with the optimal control solutions. For a comprehensive introduction to the RL methods employed here, we refer the reader to the classic textbook [sutton1998reinforcement].
V.1 Action-Value methods: Q-learning
To explain how RL works, we need to introduce the general framework. The agent can move between different states $s$ by performing actions $a$. It chooses actions $a$ randomly with a probability given by a policy function $\pi(a|s)$, which depends on the current state $s$. After each action the agent obtains some reward $r$, so the typical trajectory can be described by the sequence $s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\dots,s_{T-1},a_{T-1},r_{T},s_{T}$, where $T$ denotes the final time step (end of an episode) and $s_{T}$ is the terminal state. The sum of future rewards is called return $g_{t}=r_{t+1}+r_{t+2}+\dots+r_{T}$. The expected return, when starting in $s$ and following $\pi$ thereafter, is called the value function of state $s$
$v_{\pi}(s)=\mathbb{E}_{\pi}\left[g_{t}|s_{t}=s\right].$ (31)
The closely related concept, central for our discussion, is the action-value function $q_{\pi}(s,a)$, which determines the expected return, when the agent is in a state $s$ and chooses an action $a$
$q_{\pi}(s,a)=\mathbb{E}_{\pi}\left[g_{t}|s_{t}=s,a_{t}=a\right].$ (32)
These two objects are related to each other as $v_{\pi}(s)=\sum_{a}q_{\pi}(s,a)\pi(a|s)$.
The agent collects rewards along its way, and the goal is to find the optimal policy $\pi^{*}(a|s)$ that maximizes the expected cumulative reward over an episode. Action-value methods focus on learning the optimal function $q_{*}(s,a)$, which satisfies the Bellman equation:
[…]
In Q-learning, $q_{*}(s,a)$ is approximated iteratively through interaction with the environment using the update rule:
$label{eq:Q-update}\begin{aligned} Q(s_{t},a_{t})&\leftarrow Q(s_{t},a_{t})+\alpha\Big[\,r_{t+1}\\ &\quad+\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\,\Big],\end{aligned}$ (34)
where $Q(s,a)$ is the current approximation for the optimal action-value function and $\alpha$ is the learning rate. This update adjusts the current estimate toward a new target that combines the immediate reward and the best predicted future return. Over many iterations, provided all state-action pairs are adequately explored, $Q(s,a)$ converges to $q_{*}(s,a)$. The optimal policy is then obtained by selecting, in each state, the action that maximizes the action-value function:
$\pi^{*}(a|s)=\begin{cases}1,&\text{if $a=\operatorname*{argmax}_{a}q_{*}(s,a)$},\\ 0,&\text{otherwise}.\end{cases}$ (35)
If the transition from the equation (LABEL:eq:Bellman) to the update rule (LABEL:eq:Q-update) appears unclear, Appendix LABEL:sec:appA provides a helpful explanation of the general idea.
A key strength of Q-learning is its off-policy nature: it can learn the optimal policy even while following a different exploratory strategy. To achieve a balance between exploration and exploitation, we employ an $\varepsilon$-greedy protocol. Under this policy, the agent selects the action with the highest estimated value (the greedy action) with probability $1-\varepsilon$ and a random action with probability $\varepsilon$. A complete pseudocode for the Q-learning algorithm is presented below:
Algorithm: Q-learning
Initialize: $Q(s,a)$ arbitrarily for all states $s$ and actions $a$, except that $Q(\textit{terminal},\cdot)=0$
Loop for each episode:
Initialize starting state $s\leftarrow s_{0}$ Loop for each step of the episode: 1. Choose action $a$ from state $s$ using $\varepsilon$-greedy policy derived from $Q$ 2. Take action $a$, observe reward $r$, next state $s^{\prime}$ 3. $Q(s,a)\leftarrow Q(s,a)+\alpha\big[r+\max\limits_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\big]$ 4. $s\leftarrow s^{\prime}$ Until the episode is over Until all episodes are over
Output: $q_{*}(s,a)\approx Q(s,a)$
V.2 Policy-Gradient methods: Actor-Critic
We now turn to policy-gradient methods, a family of RL techniques that rely on directly optimizing a parameterized policy. In this approach, the agent’s behavior is described by a probability distribution $\pi(a|s,\bm{p})$, where $\bm{p}$ represents the vector of policy parameters. By smoothly adjusting the components of $\bm{p}$, the agent’s policy can be continuously modified. The central objective of such methods is to find the parameters that maximize the expected cumulative reward. This goal can be formalized as maximizing the performance function $J(\bm{p})=v_{\pi}(s_{0})$, where $s_{0}$ is the initial state. To achieve this, the parameters are updated using a gradient ascent scheme
$\bm{p}_{t+1}=\bm{p}_{t}+\alpha_{p}{\nabla_{\bm{p}}J}(\bm{p}_{t}),$ (36)
where $\alpha_{p}$ denotes the learning rate. The core difficulty lies in computing the gradient $\nabla_{\bm{p}}J(\bm{p})$, since it depends not only on the actions selected by the policy but also on the distribution of states visited during interaction with the environment.
Fortunately, this gradient can be expressed analytically thanks to the policy gradient theorem, which provides a tractable form for optimization, see Appendix LABEL:sec:appB. According to the theorem,
[…]
where $\mu(s)$ represents the stationary distribution of states under the policy $\pi$. Instead of explicitly averaging over all possible states, the same expression can be interpreted as an expectation over trajectories generated by the policy. Consequently, the gradient can be written as
$\nabla_{\bm{p}}J(\bm{p})$ $\propto$ $\mathbb{E}_{\pi}\left[\sum_{a}q_{\pi}(s_{t},a)\nabla_{\bm{p}}\pi(a|s_{t},\bm{p})\right]$
$=$ $\mathbb{E}_{\pi}\left[\sum_{a}\pi(a|s_{t},\bm{p})q_{\pi}(s_{t},a)\dfrac{\nabla_{\bm{p}}\pi(a|s_{t},\bm{p})}{\pi(a|s_{t},\bm{p})}\right].$
Next, we can replace the averaging over $a$ with the weight $\pi(a|s_{t},\bm{p})$ by sampling over $a_{t}$ generated according to the policy $\pi$, and then we can also use the identity $q_{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}[g_{t}|s_{t},a_{t}]$ to obtain
$\nabla_{\bm{p}}J(\bm{p})\propto\mathbb{E}_{\pi}\left[g_{t}\dfrac{\nabla_{\bm{p}}\pi(a_{t}|s_{t},\bm{p})}{\pi(a_{t}|s_{t},\bm{p})}\right].$ (39)
This leads to a practical stochastic gradient ascent rule of the form
$\bm{p}_{t+1}=\bm{p}_{t}+\alpha_{p}g_{t}\nabla_{\bm{p}}\ln\pi(a_{t}|s_{t},\bm{p}_{t}),$ (40)
where the proportionality constant has been absorbed into the learning rate $\alpha_{p}$. The variable $g_{t}$ represents the total return accumulated from time $t$ onward, implying that parameter updates can only be performed after completing an entire episode. This approach is known as the REINFORCE algorithm [williams1992simple].
A notable improvement to REINFORCE involves introducing a baseline term $b(s)$, which helps reduce the variance of the gradient estimate without introducing bias. This modification is justified by the observation that
$\sum_{a}b(s)\nabla_{\bm{p}}\pi(a|s,\bm{p})=b(s)\nabla_{\bm{p}}\sum_{a}\pi(a|s,\bm{p})=0,$ (41)
since $\sum_{a}\pi(a|s,\bm{p})=1$. Hence, one can replace $q_{\pi}(s,a)$ with $q_{\pi}(s,a)-b(s)$ in the policy gradient theorem (LABEL:eq:theorem), leading to the modified update rule
$\bm{p}_{t+1}=\bm{p}_{t}+\alpha_{p}(g_{t}-b(s_{t}))\nabla_{\bm{p}}\ln\pi(a_{t}|s_{t},\bm{p}_{t}).$ (42)
This modification leaves the expected value of the gradient unchanged but may drastically reduce its variance. In practice, an effective choice for the baseline is the state value function itself, $b(s_{t})\approx v_{\pi}(s_{t})$.
Building upon this idea, the one-step actor-critic algorithm introduces an even more efficient mechanism by replacing $g_{t}$ with a one-step return. Instead of waiting for the end of an episode, the algorithm uses the estimate $g_{t}\approx r_{t+1}+v_{\pi}(s_{t+1})$ to compute an immediate update. To make this feasible, the value function is approximated by a differentiable function $V(s,\bm{w})$ parameterized by a weight vector $\bm{w}$. The parameters $\bm{w}$ are trained to minimize the mean-squared error between the true value function and its approximation $v_{\pi}(s_{t})-V(s_{t},\bm{w})]^{2}$ using a stochastic semi-gradient method. The corresponding update rule is
$\bm{w}_{t+1}=\bm{w}_{t}-\dfrac{\alpha_{w}}{2}\nabla_{\bm{w}}[v_{\pi}(s_{t})-V(s_{t},\bm{w}_{t})]^{2}$ (43)
$=\bm{w}_{t}+\alpha_{w}[v_{\pi}(s_{t})-V(s_{t},\bm{w}_{t})]\nabla_{\bm{w}}V(s_{t},\bm{w}_{t}).$
Since the true value $v_{\pi}(s_{t})$ is unknown, it is replaced with an unbiased estimate based on the observed reward and the value of the next state, $v_{\pi}(s_{t})\approx r_{t+1}+V(s_{t+1},\bm{w}_{t})$. Strictly speaking, this estimate depends on $\bm{w}_{t}$, but we ignore this when computing the gradient in the update rule, and for this reason the method is called semi-gradient.
With these definitions, both the policy and the value function parameters can be updated after each interaction step. The resulting learning rules are
$\bm{p}_{t+1}=\bm{p}_{t}$ $+$ $\alpha_{p}\Big[r_{t+1}+V(s_{t+1},\bm{w}_{t})$
- $V(s_{t},\bm{w}_{t})\Big]\nabla_{\bm{p}}\ln\pi(a_{t}|s_{t},\bm{p}_{t}),$
$\bm{w}_{t+1}=\bm{w}_{t}$ $+$ $\alpha_{w}\Big[r_{t+1}+V(s_{t+1},\bm{w}_{t})$
- $V(s_{t},\bm{w}_{t})\Big]\nabla_{\bm{w}}V(s_{t},\bm{w}_{t}).$
In this formulation, the first equation adjusts the actor, i.e., the policy $\pi(a|s,\bm{p})$ responsible for choosing actions, while the second equation updates the critic, i.e., the value function $V(s,\bm{w})$ that evaluates them. The actor learns to favor actions that produce higher-than-expected rewards, guided by the error $r_{t+1}+V(s_{t+1},\bm{w}_{t})-V(s_{t},\bm{w}_{t})$, while the critic improves its estimate of expected returns using the same signal. The critic “evaluates” the actor’s choices, and the actor uses this feedback to refine its policy, reinforcing beneficial actions and discouraging poor ones. A complete pseudocode for the one-step actor-critic algorithm is presented below:
Algorithm: One-Step Actor-Critic
Initialize: Differentiable $\pi(a|s,\bm{p})$ and $V(s,\bm{w})$ with some initial parameters $\bm{p}$ and $\bm{w}$, $V(terminal,\cdot)=0$
Loop for each episode:
Initialize starting state $s\leftarrow s_{0}$ Loop for each step of the episode: 1. Choose action $a$ from state $s$ using $\pi(\cdot|s,\bm{p})$ 2. Take action $a$, observe reward $r$, next state $s^{\prime}$ 3. $\varkappa\leftarrow r+V(s^{\prime},\bm{w})-V(s,\bm{w})$
4. $\bm{w}\leftarrow\bm{w}+\alpha_{w}\varkappa\nabla_{\bm{w}}V(s,\bm{w})$
$\bm{p}\leftarrow\bm{p}+\alpha_{p}\varkappa\nabla_{\bm{p}}\ln\pi(a|s,\bm{p})$
5. $s\leftarrow s^{\prime}$
Until the episode is over
Until all episodes are over
Output: $\pi^{*}(a|s)\approx\pi(a|s,\bm{p})$
Note that once training is complete, we evaluate the learned control in a testing phase by selecting, at each state $s_{t}$, the greedy action $a_{t}$ with the highest probability under the learned policy, i.e. $a_{t}=\operatorname*{argmax}_{a}\pi(a|s_{t},\bm{p})$. This deterministic choice allows us to highlight the most likely trajectories implied by the policy, making it easier to visualize and interpret the agent’s characteristic behavior.
[…]
V.3 States, Actions, Rewards
Next, we outline how the general RL framework introduced above can be applied to Zermelo’s navigation problem. The definitions of states, actions and rewards follow Ref. [biferale2019zermelo], with only minor modifications in our implementation.
The state space is constructed by discretizing the spatial domain $L\times L$ into square tiles or cells $s_{i}$, with $i=1,2,\dots,N_{s}$, each of size $\delta\times\delta$. We set $\delta=L/N_{tile}$, so that in total we have $N_{s}=N_{tile}^{2}$ states. Roughly saying, this coarse graining implies that the agent resolves its position only up to the cell size $\delta$. The agent is allowed to change its steering direction only at discrete time intervals of duration $\Delta t=\delta/(2u_{0})$, where $u_{0}$ is taken to be of the order of the maximum flow velocity. The action space is also discretized: the control variable $\theta_{t}$ can take one of $N_{a}=8$ equally spaced angles $\theta_{j}=-\pi+2\pi j/N_{a}$ with $j=1,2,\dots,N_{a}$. During each time step of duration $\Delta t$, the agent’s trajectory is advanced via numerical integration of the governing equation of motion (LABEL:eq:motion) with the chosen steering angle held fixed.
We define one episode as a single attempt to reach the target region. An episode terminates either when the agent enters a circle of radius $d_{B}$ centered at the target position $\bm{r}_{B}$, or when a maximum allowed time $T_{max}$ has elapsed. Episodes also terminate prematurely if the agent collides with the domain boundary. Boundary collisions incur an additional negative reward of magnitude $2T_{max}$, providing a strong penalty that discourages such events. The reward structure is designed to favor short travel times. At each action step, the immediate reward is
$r_{t}=-\Delta t+\dfrac{|\bm{r}_{B}-\bm{R}_{t-\Delta t}|}{V_{0}}-\dfrac{|\bm{r}_{B}-\bm{R}_{t}|}{V_{0}},$ (46)
where $V_{0}$ is the slip velocity of the agent and $\bm{R}_{t}$ denotes its position at time $t$. The first term imposes a constant running cost proportional to time spent, ensuring that long trajectories accumulate a large negative reward. The second and third terms measure the improvement in a distance-based potential and is known to preserve optimality while significantly accelerating convergence during training [biferale2019zermelo, gunnarson2021learning]. When summed over a successfully completed episode (the agent has reached the target), these terms result in a nearly constant addition $\approx|\bm{r}_{B}-\bm{r}_{A}|/V_{0}$. To enhance robustness with respect to small perturbations of the initial condition, each episode is initialized by sampling the agent’s starting point $\bm{R}_{0}$ uniformly within a circle of radius $d_{A}$ centered at the nominal initial position $\bm{r}_{A}$. Throughout this work, we set $d_{A}=d_{B}\approx\delta\ll|\bm{r}_{B}-\bm{r}_{A}|$.
In the above formulation, the system state is characterized solely by the agent’s spatial position. Although this assumption is appropriate for steady background flows, it may complicates learning or even become restrictive when the flow depends explicitly on time. In such cases, the optimal action at a fixed spatial location can vary as the flow evolves, and a state representation that excludes temporal information may hinder the algorithm’s ability to identify optimal trajectories – especially those exhibiting complex features such as self-intersections. Nevertheless, for the time-dependent scenarios examined in this work, we found that RL algorithms employing this reduced state representation can still cope with temporal variability and successfully learn effective control strategies.
[…]
Given the discrete state-action space, with $i=1,2,\dots,N_{s}$ indexing the states and $j=1,2,\dots,N_{a}$ indexing the actions, the action-value function $Q(s,a)$ used in the Q-learning algorithm can be represented explicitly as a matrix $Q_{ij}$ of dimension $N_{s}\times N_{a}$. Each entry $Q_{ij}$ corresponds to the estimated value of taking action $a_{j}$ in state $s_{i}$. The policy-gradient methods require specifying a parametric form for the policy $\pi(a|s,\bm{p})$, where $\bm{p}$ denotes the set of trainable parameters. A convenient and widely used choice is the softmax parameterization:
$\pi(a_{j}|s_{i},\bm{p})=\dfrac{\exp(p_{ij})}{\sum_{k=1}^{N_{a}}\exp(p_{ik})},$ (47)
where $p_{ij}$ is the $(i,j)$-th element of an $N_{s}\times N_{a}$ parameter matrix. Thus, each state-action pair $(s_{i},a_{j})$ is associated with a scalar parameter $p_{ij}$ that determines the probability of selecting action $a_{j}$ when the agent is in state $s_{i}$. For the state-value function, we also choose one of the most common parameterizations,
$V(s_{i},\bm{w})=w_{i},$ (48)
where $w_{i}$ is the $i$-th component of a parameter vector $\bm{w}$ of length $N_{s}$. This representation assigns one free parameter to each state, keeping the function approximation simple. With these choices, the update rules (V.2) and (V.2) for a one-step actor-critic method take the form:
$\varkappa\leftarrow r+w[s_{t+1}]-w[s_{t}],$ (49)
$w[s_{t}]\leftarrow w[s_{t}]+\alpha_{w}\varkappa,$ (50)
$p[s_{t},a_{t}]\leftarrow p[s_{t},a_{t}]+\alpha_{p}\varkappa\,(1-\pi(a_{t}|s_{t})),$ (51)
$p[s_{t},b\neq a_{t}]\leftarrow p[s_{t},b\neq a_{t}]-\alpha_{p}\varkappa\,\pi(b\neq a_{t}|s_{t}).$ (52)
V.4 Example 1: Steady vortex flow
First, we apply the RL algorithms to the navigation task in the steady vortex flow, see Section LABEL:sec:Ex_A. The flow domain is restricted to the square $0.3,1.2]^{2}$, which is discretized into $N_{s}=25^{2}$ uniform cells defining the states. The uncertainties in the initial and target points are controlled by $d_{A}=d_{B}=0.04$. Each episode has a maximum duration of $T_{max}=5$. The agents are trained for $n_{train}=10^{5}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we use a constant learning rate $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ after each episode. In the one-step actor-critic algorithm, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.04$ and $\alpha_{w}=0.08$, respectively.
The evolution of the total reward $r_{tot}=\sum_{t}r_{t}$ during training, averaged over a sliding window of $250$ episodes, is shown for both algorithms in Fig. LABEL:fig:10a. In each case, the total reward eventually stabilizes, indicating that the learning dynamics have reached a stationary regime and the algorithms have converged. Examples of trajectories obtained from the learned policies are presented in Fig. LABEL:fig:10b. For each method, we display ten realizations initialized uniformly at random within a disk of radius $d_{A}$ centered at $\bm{r}_{A}$. For reference, the optimal trajectory computed in Section LABEL:sec:OPC_A is shown alongside the RL-generated paths. The trajectories produced by Q-learning closely follow the optimal control solution; the small deviations that remain can be attributed primarily to the discretization inherent in our RL formulation. In contrast, the one-step actor-critic method yields trajectories that exhibit systematic deviations from the optimal path. These deviations are robust: they persist across independent training runs and remain even after varying the training hyperparameters. Apparently, this behavior can be related to the on-policy nature of one-step actor-critic algorithm, which may converge to local minimum of the policy-gradient landscape.
[…]
The performance difference between the two methods is illustrated by the arrival-time statistics shown in Fig. LABEL:fig:10c. All evaluated trajectories reach the target, implying that both methods achieve a zero failure rate. However, the Q-learning agent achieves a slightly shorter mean travel time, $\langle T\rangle\approx 2.14$, compared with $\langle T\rangle\approx 2.18$ for the one-step actor-critic agent. Therefore, while both RL approaches successfully solve the navigation task, Q-learning provides a closer approximation to the optimal control solution and delivers better performance.
V.5 Example 2: Steady sink and time-varying vortex flow
Next, we apply the Q-learning and one-step actor-critic algorithms to the navigation task in the steady sink and time-dependent vortex flow, see Section LABEL:sec:Ex_B. The training hyperparameters are identical to those used in the previous example, except that for the one-step actor-critic method we slightly increase the learning rates to $\alpha_{p}=0.05$ for the actor and $\alpha_{w}=0.1$ for the critic. The corresponding results are summarized in Fig. LABEL:fig:11.
Both algorithms converge successfully in this time-dependent external flow, as illustrated in Fig. LABEL:fig:11a, despite the fact that the state representation relies solely on the agent’s position. The trajectories produced by the learned policies are reasonably close to the optimal control solution derived in Section LABEL:sec:OPC_B, see Fig. LABEL:fig:11b. A visualization of the full time-dependent evolution is available in the videos provided in Ref. [repository]. All evaluated trajectories reach the target, demonstrating a zero failure rate for both approaches. The arrival-time statistics in Fig. LABEL:fig:11c indicate that the two algorithms perform comparably, with mean travel times of approximately $\langle T\rangle\approx 1.065$.
V.6 Example 3: Taylor-Green vortices
The next example concerns the Taylor-Green flow described in Section LABEL:sec:Ex_C. The flow domain $0,2\pi]^{2}$ is discretized into $N_{s}=30^{2}$ uniform square cells that define the state space. The uncertainties in the initial and target positions are set to $d_{A}=d_{B}=0.2$. The maximum episode duration is given by $T_{\max}=2|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}=0.1$ denotes the agent’s self-propulsion speed. The agents are trained for $n_{train}=2\cdot 10^{4}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we employ a constant learning rate of $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9995$ after each episode. For the one-step actor-critic method, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
Figure LABEL:fig:12a shows the learning curves obtained for the Q-learning and one-step actor-critic algorithms. In both cases, the curves level off at relatively high total rewards, indicating that each method successfully learns an effective control policy. Representative trajectories generated by the learned policies are displayed in Fig. LABEL:fig:12b, alongside the optimal path computed in Section LABEL:sec:OPC_C. As illustrated, the Q-learning agent typically proceeds directly toward the target along a route close to the optimal one, although it may complete a single revolution around a nearby vortex when its initial state makes immediate escape impossible. The actor-critic agent performs less reliably: in addition to this initial revolution, it is sometimes captured by a second vortex, which forces an additional turn.
These observations are further supported by the arrival-time statistics shown in Fig. LABEL:fig:12c. The first peak corresponds to near-optimal trajectories; the second arises from trajectories that include one extra loop within a Taylor-Green vortex; and the third reflects trajectories with two such additional rotations. Because of uncertainty in the initial position and the larger size of the target region, some trajectories terminate faster than the optimal-control prediction (LABEL:eq:opt-control-taylor-green). Note that all tested trajectories for both RL algorithms arrive at the target, yielding a zero failure rate. The appearance of the third peak leads to a slightly longer mean travel time for the actor-critic agent, $\langle T\rangle\approx 13.1$, compared with $\langle T\rangle\approx 12.9$ for the Q-learning agent. However, this difference is only a few percent, and overall both algorithms solve the task successfully.
For comparison, in Ref. [repository] we have also implemented a trivial policy, where the steering angle selected by the agent at each $\Delta t$ is given by the action that points most directly toward the target among the $N_{a}=8$ different possibilities. In this case, the agent reaches the target only with a probability of $\approx 0.3$ (depending on its random starting position), and in the remaining attempts it ends up trapped in Taylor-Green vortices, from which this naive strategy does not allow to escape.
[…]
V.7 Example 4: Snapshot of 2D Turbulence
[…]
sec:RL_2d_turb_snap
Next, we apply the RL algorithms to a navigation task in the steady velocity field obtained from a snapshot of two-dimensional turbulence, see Section LABEL:sec:Ex_D. The flow domain, $0,2\pi]^{2}$, is discretized into $N_{s}=30^{2}$ uniform square cells, which serve as the state space. Uncertainty in both the initial and target positions is characterized by $d_{A}=d_{B}=0.2$. The maximum episode duration is set to $T_{\max}=5|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}$ is the agent’s self-propulsion speed. Agents are trained for $n_{train}=10^{5}$ episodes and evaluated over an additional $n_{eval}=5\times 10^{3}$ episodes. For Q-learning, we use a constant learning rate $\alpha=0.04$ and gradually reduce the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ at the end of each episode. For the one-step actor-critic algorithm, the actor and critic learning rates are kept constant at $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
Figure LABEL:fig:13a shows the evolution of the total reward throughout training. For both RL algorithms, the learning curves level off, indicating convergence. Representative trajectories generated by the learned policies are displayed in Fig. LABEL:fig:13b, together with the optimal control trajectory from Section LABEL:sec:OPC_D. As seen in the figure, the actor-critic agent’s trajectories closely track the optimal control path, whereas the Q-learning trajectories exhibit small deviations in the latter part of the journey. Consistently, the actor-critic agent achieves a shorter average travel time, $\langle T\rangle\approx 0.87$, compared with $\langle T\rangle\approx 0.96$ for the Q-learning agent, see Fig. LABEL:fig:13c.
This difference does not contradict the observation in Fig. LABEL:fig:13a, where the Q-learning agent attains a higher total reward at the end of training. The discrepancy arises because, during evaluation, the actor-critic agent selects the most probable actions from the learned policy, which yield higher rewards and consequently shorter travel times. Overall, all evaluated trajectories successfully reach the target, demonstrating a zero failure rate for both algorithms. Nonetheless, in the considered example the one-step actor-critic method provides a closer approximation to the optimal control solution and delivers better performance. Note also that the trivial policy implemented in Ref. [repository] completely fails to solve the considered navigation task, yielding a failure rate of $1$. Under this strategy, the agent eventually enters regions where the external flow cancels its self-propulsion, causing it to stall. As a result, it becomes effectively immobilized and fails to reach the target.
[…]
V.8 Example 5: 2D Turbulence
[…]
sec:RL_2d_turb_time
In our final example, we consider navigation in a time-dependent 2d turbulent flow, see Section LABEL:sec:Ex_D. The Q-learning hyperparameters are identical to those used for the steady 2d turbulence snapshot, while for the one-step actor-critic algorithm we increase the learning rates to $\alpha_{p}=0.04$ for the actor and $\alpha_{w}=0.08$ for the critic. The obtained results are summarized in Fig. LABEL:fig:14.
Both methods converge, as shown by the learning curves in Fig. LABEL:fig:14a. The trajectories generated by their learned policies are similar to each other but diverge from the optimal-control solution presented in Section LABEL:sec:OPC_E, see Fig. LABEL:fig:14b. The discrepancies arise mainly in the second half of the path: the optimal control solution identifies a shorter and faster route through a region of weak external flow that the agent can overcome, whereas the RL policies exploit a strong background current pointing toward the goal. A corresponding video demonstration is provided in Ref. [repository]. Evidently, the shorter path requires more coordinated actions and is therefore more difficult for the RL agents to discover. The longer RL-derived path is more stable and still yields a high total reward, so both algorithms converge to it.
The performance difference between Q-learning and the actor-critic method is reflected in the arrival-time statistics in Fig. LABEL:fig:14c. All evaluated trajectories reach the target, indicating zero failure rate for both algorithms. However, the actor-critic agent exhibits a higher mean arrival time and a larger variance, implying that the Q-learning solution is slightly more efficient and robust. Overall, both methods successfully complete the navigation task in a complex 2d turbulent flow and demonstrate comparable performance.
Finally, Ref. [repository] also includes a trivial policy for this scenario. It fails to reach the target in most runs, with a failure rate close to $0.95$. Even when it succeeds, the travel time is very long and the resulting trajectories are not optimal.
VI Incomplete Flow Information
[…]
sec:6
So far, we have seen that RL algorithms successfully solve navigation problems in various background flows with zero failure rate and exhibit strong robustness to perturbations in the agent’s initial position. These results suggest that RL methods may also remain stable under small fluctuations in the external velocity field that may arise, among other reasons, from incomplete information about the surrounding flow. This aspect is particularly crucial for predicting movement in real atmospheric or oceanic environments, where small-scale turbulent fluctuations are difficult to forecast. In this section, we examine this issue in detail.
VI.1 Methodology
In two-dimensional turbulent flows, most of the kinetic energy is concentrated at large spatial scales, see Fig. LABEL:fig:1. In the inverse energy cascade range, $k<k_{f}$, where $k_{f}$ denotes the forcing wavenumber, the energy spectrum follows the scaling $E(k)\propto k^{-5/3}$. Consequently, the characteristic velocity associated with motions of size $r$ can be estimated as $v_{r}\sim(\epsilon r)^{1/3}$, where $\epsilon$ is the energy flux in the inverse cascade or, equivalently, the power injected into the system [boffetta2012two]. The characteristic timescale for velocity variations at scale $r$ then becomes $t_{r}\sim r/v_{r}\sim\epsilon^{-1/3}r^{2/3}$. Thus, the energetically dominant large scales evolve slowly and are relatively easy to predict, whereas small-scale fluctuations change rapidly and are inherently difficult to forecast with precision.
In what follows, we assume that the agent has access to an accurate prediction of all flow modes with wavenumbers $k\leq k_{max}$, while possessing no information about the smaller-scale fluctuations. Practically, this means that we take the velocity field produced by the DNS and remove all Fourier modes with $k>k_{max}$. By varying the filtering threshold $k_{max}$, we can directly control the amount of flow information available to the agent.
We then train the agent using RL algorithms on this coarse-grained flow field and determine the resulting navigation strategies. Finally, we evaluate these strategies in the full velocity field to assess whether the agent can still successfully reach its target when exposed to the previously unseen small-scale fluctuations. Below, we examine two representative navigation scenarios: one in a steady snapshot of the 2d turbulent flow, and another in the fully time-dependent 2d turbulent field. In our examples, the forcing scale is fixed at $k_{f}=5.5$, and we progressively reduce the filtering scale, considering $k_{max}=\{10,5,3,2\}$.
[…]
VI.2 Example 1: Snapshot of 2D Turbulence
We begin by examining the navigation problem in a single snapshot of the 2d turbulent flow. The training hyperparameters for both the Q-learning and one-step actor-critic algorithms are identical to those used in Section LABEL:sec:RL_2d_turb_snap. For all considered filtering thresholds, $k_{max}=\{10,5,3,2\}$, the learning curves for both methods reach a clear plateau, indicating convergence toward effective control policies.
Representative trajectories obtained from these learned policies – evaluated in the full (unfiltered) velocity field – are shown in Fig. LABEL:fig:15. For comparison, we also display the optimal control trajectory computed in Section LABEL:sec:OPC_D. As illustrated, the learned policies remain robust even when the agent is provided only with coarse-grained flow information: for $k_{max}\geq 3$, both algorithms reliably guide the agent to the target along paths that remain reasonably close to the optimal solution. Consistent with our earlier findings, the one-step actor-critic agent performs slightly better than the Q-learning agent. However, the difference in their mean travel times does not exceed $10\%$, and overall both RL methods succeed in solving the task with zero failure rate, see Table LABEL:tab:1.
When the filtering becomes more severe, at $k_{max}=2$, the situation changes qualitatively. The Q-learning agent is no longer able to reach the target at all. The actor-critic agent performs somewhat better, achieving success in roughly $20\%$ of trials, but its average travel time nearly doubles compared with the optimal control trajectory. In summary, removing velocity modes with wavenumbers $k>2$ significantly distorts the flow structure, to the point where the learned policies become practically useless.
Table $\labeltab:1$Example of the 2d turbulence snapshot: Mean travel time $\langle T\rangle$ of successful target-reaching attempts and failure rates across different filtering thresholds $k_{max}$.
Q-learning One-Step Actor-Critic
$k_{max}$ $\langle T\rangle$ failure rate $\langle T\rangle$ failure rate
$\infty$ $0.96$ 0 $0.87$ 0
$10$ $0.93$ 0 $0.87$ 0
55 $0.94$ 0 $0.91$ 0
33 $0.91$ 0 $0.88$ 0
22 11 $1.65$ $0.77$
VI.3 Example 2: 2D Turbulence
The next example focuses on a navigation problem in time-dependent 2d turbulent flow. The training hyperparameters for both the Q-learning and one-step actor-critic methods are selected to be identical to those used earlier in Section LABEL:sec:RL_2d_turb_time. For all filtering thresholds $k_{max}=\{10,5,3,2\}$, we achieved convergence of the RL algorithms toward effective control policies, as confirmed by the learning curves reaching a plateau.
Figure LABEL:fig:16 shows the agent trajectories corresponding to the learned policies for different filtering scales $k_{max}$ of velocity field. For reference, we also display the trajectory obtained from the optimal control theory in Section LABEL:sec:OPC_E. Consistent with our earlier findings, the RL algorithms converge to a trajectory distinct from the optimal control path. This RL-derived solution remains robust with respect to the removal of small-scale velocity fluctuations with $k\gtrsim k_{f}$. When the filtering scale $k_{max}$ is reduced further, the agent trajectories become increasingly irregular. Moreover, in rare cases, the agent fails to reach the target, and for those attempts that do succeed, the average travel time increases, see Table LABEL:tab:2. When comparing the Q-learning and one-step actor-critic methods, the action-value approach seems to performs slightly better at low $k_{max}$.
At a qualitative level, the example of time-dependent 2d turbulent flow considered here behaves similarly to the previously studied example of 2d turbulent snapshot. The trajectories produced by the RL algorithms exhibit robustness to the absence of small-scale velocity field information; however, once the filtering scale approaches the characteristic size of the flow structures that guide the agent’s path selection, the performance of the methods deteriorates noticeably. For 2d turbulence, the critical filtering scale $k_{max}$ lies within the inverse energy cascade range.
Table $\labeltab:2$Example of the time-dependent 2d turbulent flow: Mean travel time $\langle T\rangle$ of successful target-reaching attempts and failure rates across different filtering thresholds $k_{max}$.
Q-learning One-Step Actor-Critic
$k_{max}$ $\langle T\rangle$ failure rate $\langle T\rangle$ failure rate
$\infty$ $0.94$ 0 $0.97$ 0
$10$ $0.93$ 0 $0.94$ 0
55 $0.99$ 0 $0.97$ 0
33 $0.99$ $0.05$ $1.27$ $0.01$
22 $1.34$ $0.04$ $3.01$ $0.27$
[…]
VII Conclusion
[…]
sec:7
In this work, we conducted a systematic investigation of optimal navigation in two-dimensional flows of increasing dynamical complexity, ranging from steady vortical fields to fully developed turbulence in the inverse energy cascade regime. Using classical optimal control theory, we demonstrated that time-minimizing trajectories can be computed accurately even in complex environments; however, their practical applicability becomes limited in complex flows due to strong sensitivity to initial conditions and the instability of the resulting optimal paths. These issues highlight the fundamental challenges associated with implementing pre-computed optimal-control strategies in realistic, chaotic settings where small perturbations and incomplete knowledge of the environment are unavoidable.
To address these limitations, we employed RL as a model-free alternative capable of adapting to complex flow structures without relying on explicit knowledge of the governing dynamics. By comparing two canonical classes of RL algorithms – action-value (Q-learning) and policy gradient (one-step actor-critic) methods – we showed that both approaches reliably produce robust navigation strategies across all considered flow configurations. While the RL-derived trajectories do not always coincide with the optimal-control solutions, they remain stable and consistently achieve successful target acquisition, even in fully time-dependent turbulence.
A key result of our study is the strong generalization capability of RL agents trained on coarse-grained velocity fields. Agents exposed only to large-scale flow information were nevertheless able to navigate efficiently within the full (unfiltered) turbulent environment. This robustness to missing small-scale data is particularly promising for real-world applications, where environmental forecasts are typically low-resolution and subject to uncertainty.
The framework introduced here also offers a flexible platform for future studies. The use of deep RL, where action-value functions and policies are represented by neural networks, would eliminate the need for state-action discretization and enable navigation in more complex geometries or higher-dimensional velocity fields. Next, the objective functional itself can be generalized: besides minimizing travel time, one may incorporate additional criteria such as energy expenditure, fuel consumption, or multi-objective trade-offs between efficiency and safety [buzzicotti2020optimal]. Another important direction of research can be related to more realistic agent models. Incorporating inertia, turning-rate limitations and stochastic forcing representing environmental noise would further bridge the gap between theory and practical deployment [colabrese2017flow, alageshan2020machine, piro2022optimal].
From a methodological viewpoint, the coexistence of ground truth optimal-control solutions and RL-based strategies within the same testbed creates a rare opportunity. It allows quantitative comparison between exact minimum-time trajectories and learned policies in data-driven control. Because the navigation problem is physically transparent, the framework is also well suited for use in teaching, providing an accessible setting in which students can explore nonlinear dynamics, turbulent transport, and modern machine-learning techniques within a unified context.
Overall, our findings demonstrate that RL provides a powerful and resilient framework for navigation in complex flow environments, complementing classical optimal control theory and overcoming its primary limitations in chaotic regimes. These results open the door to developing practical autonomous navigation strategies for micro-robots, underwater vehicles, and other active agents operating in realistic oceanic and atmospheric flows.
Acknowledgements.
The work was supported by the Ministry of Science and Higher Education of the Russian Federation and by the Basic research program of HSE University.
Appendix A Incremental update
[…]
sec:appA
Assume for a moment that we are interested in estimating the mean value of a random number generator. At each step, we obtain a single sample, and just before the $n$-th step, our best estimate of the mean is given by
[…]
From this definition, one can derive a recursive update rule that allows computing the next estimate without explicitly storing all past samples:
[…]
where $Q_{1}=0$. The main advantage of this incremental form is its computational efficiency: instead of maintaining the entire history of samples, we only need the current estimate $Q_{n}$ and the most recent observation $r_{n}$. Each new sample slightly adjusts the estimate in the direction of the new data point, with the step-size determined by $1/n$.
More generally, this recursive formulation can be expressed in the abstract form
$\textit{New Estimate}\leftarrow$ Old Estimate $\textit${Old Estimate }+ (55)
$\alpha\left(\textit{Target}-\textit{Old Estimate}\right),$
where the target represents a potentially noisy indicator of the true value toward which the estimate should move, and $\alpha$ is the step-size or learning rate that controls how strongly the estimate is updated. In the present case, the target corresponds to the new sample $r_{n}$, and the learning rate is $\alpha=1/n$. Because this learning rate decays over time, each update becomes smaller, which guarantees convergence of $Q_{n}$ to the true mean as $n\to\infty$ in accordance with the law of large numbers, see Eq. (LABEL:eq:A1).
The choice of learning rate schedule determines how the estimate responds to new data. When a decaying learning rate $\alpha=1/n$ is used, all samples contribute equally to the final estimate. This approach is suitable when the underlying random process is stationary. If, on the other hand, the process is non-stationary and its statistical properties vary with time, assigning equal weight to all past samples may be sub-optimal. In such cases, it is preferable to emphasize more recent observations by using a constant learning rate $\alpha=\text{const}$. With a constant learning rate, the estimate can be written as
$Q_{n+1}=\sum_{i=1}^{n}\alpha(1-\alpha)^{n-i}r_{i},$ (56)
which corresponds to an exponentially weighted moving average where more recent samples have greater influence. This formulation enables the estimate to adapt to changing conditions, but it also introduces persistent fluctuations around the true mean due to the most recent samples. The magnitude of these fluctuations decreases as the learning rate becomes smaller, at the cost of slower adaptation to new information.
In practice, choosing an appropriate learning rate or decay schedule involves balancing adaptability and stability. If the learning rate decays too quickly, the estimate may become insensitive to new data; if it decays too slowly, convergence to the true mean may not occur. Modern machine learning algorithms address this trade-off through adaptive learning rate methods, such as Adam [kingma2014adam], RMSProp [tieleman2012lecture], or Adagrad [duchi2011adaptive], which automatically adjust the effective step-size based on the observed variance of updates, see also Ref. [mehta2019high]. These methods generalize the simple recursive rule presented above and form the foundation of many contemporary optimization techniques.
Appendix B Policy Gradient Theorem
[…]
sec:appB
The proof of the policy gradient theorem presented here follows closely the derivation given in Ref. [sutton1998reinforcement], and is included for completeness. Throughout the derivation, we assume that the policy $\pi(a|s)$ depends on a parameter vector $\bm{p}$, and all gradients are taken with respect to these parameters. For notation compactness, this dependence is left implicit.
For any state $s$, the state-value function can be expressed in terms of the action-value function as $v_{\pi}(s)=\sum_{a}q_{\pi}(s,a)\pi(a|s)$. Taking the gradient of this expression yields
$label{eq:b1}\nabla v_{\pi}(s)=\sum_{a}\left[\nabla\pi(a|s)q_{\pi}(s,a)+\pi(a|s)\nabla q_{\pi}(s,a)\right].$ (57)
To evaluate the second term, we represent the action-value function in terms of the environment’s transition dynamics. Given the joint probability $p(s^{\prime},r|s,a)$ of reaching state $s^{\prime}$ and receiving reward $r$ after taking action $a$ in state $s$, the function $q_{\pi}(s,a)$ can be written as
$q_{\pi}(s,a)=\sum_{s^{\prime},r}p(s^{\prime},r|s,a)(r+v_{\pi}(s^{\prime})).$ (58)
Since the environment’s transition model does not depend on the policy parameters, differentiation gives
$label{eq:b3}\nabla q_{\pi}(s,a)=\sum_{s^{\prime}}p(s^{\prime}|s,a)\nabla v_{\pi}(s^{\prime}),$ (59)
where $p(s^{\prime}|s,a)=\sum_{r}p(s^{\prime},r|s,a)$ denotes the marginal transition probability. Substituting Eq. (LABEL:eq:b3) into Eq. (LABEL:eq:b1) leads to a recursive form,
[…]
The term involving $\nabla v_{\pi}(s^{\prime})$ can itself be expanded using Eq. (LABEL:eq:b4), and this substitution can be applied recursively to capture contributions from all future states along a trajectory generated by the policy. After the first substitution, we obtain
$\nabla v_{\pi}(s)=\sum_{a}\Big[\nabla\pi(a|s)q_{\pi}(s,a)$
$+\pi(a|s)\sum_{s^{\prime}}p(s^{\prime}|s,a)\sum_{a^{\prime}}\Big[\nabla\pi(a^{\prime}|s^{\prime})q_{\pi}(s^{\prime},a^{\prime})$ (61)
$+\pi(a^{\prime}|s^{\prime})\sum_{s^{\prime\prime}}p(s^{\prime\prime}|s^{\prime},a^{\prime})\nabla v_{\pi}(s^{\prime\prime})\Big]\Big].$
Carrying out this expansion indefinitely yields an infinite series representing the influence of all possible state transitions:
$\nabla v_{\pi}(s)=\sum_{x}\sum_{k=0}^{\infty}P(s\to x,k,\pi)\sum_{a}\nabla\pi(a|x)q_{\pi}(x,a),$ (62)
where $P(s\to x,k,\pi)$ denotes the probability of reaching state $x$ from $s$ in exactly $k$ steps under policy $\pi$. For instance, $P(s\to x,1,\pi)=\sum_{a}p(x|s,a)\pi(a|s)$.
The overall performance objective of the policy is defined as $J(\bm{p})=v_{\pi}(s_{0})$, where $s_{0}$ is the initial state (or drawn from an initial distribution). Taking the gradient of this quantity and substituting the above expansion, we obtain
$\nabla J(\bm{p})=\sum_{s}\left(\sum_{k=0}^{\infty}P(s_{0}\to s,k,\pi)\right)\sum_{a}\nabla\pi(a|s)q_{\pi}(s,a).$ (63)
The inner summation represents the expected visitation frequency of state $s$, which is proportional to the stationary distribution of states $\mu(s)$ under policy $\pi$. Using this notation, the gradient of the performance objective can be written in the form
$\nabla J(\bm{p})\propto\sum_{s}\mu(s)\sum_{a}\nabla\pi(a|s)q_{\pi}(s,a).$ (64)